OpenStudy (anonymous):

I've got the complex number: \[z_{0}=2e^{\dfrac{i\pi}{6}}\]and i can't figure out Im(z) and Re(z]

OpenStudy (anonymous):

the e^ notation has a circle and you are moving on it. \[\frac{ \Pi }{ 6 }\]

OpenStudy (anonymous):

there's an equation are you allowed to use it? \[\cos x + i \sin x\]

OpenStudy (anonymous):

Yes i guess so, i have to find the polar coordinates

OpenStudy (anonymous):

i sin x is the complex part of the complex pointer in the polar graph and x = pi/6

OpenStudy (anonymous):

is the imaginary part, I mean

OpenStudy (anonymous):

for the number you got, it can have the extreme cases Re=0, Im=2 or Re=2, Im=0 the distribution depends on the angle pi/6 and the triangle's sides that are opened by the complex pointer

OpenStudy (anonymous):

in other words: Im(z) =2* sin pi/6 and Re(z) =2* cos pi/6

OpenStudy (anonymous):

Thank you very much for your help, i'll be back in a min, i'll just have to make dinner :)